# Optimal Execution with Reinforcement Learning

### v1

@@ Abstract
This study investigates the development of an optimal execution strategy through reinforcement learning, aiming to determine the most effective approach for traders to buy and sell inventory within a limited time frame. Our proposed model leverages input features derived from the current state of the limit order book.
To simulate this environment and overcome the limitations associated with relying on historical data, we utilize the multi-agent market simulator ABIDES, which provides a diverse range of depth levels within the limit order book.
We present a custom MDP formulation followed by the results of our methodology and benchmark the performance against standard execution strategies. Our findings suggest that the reinforcement learning-based approach demonstrates significant potential.
Keywords: Optimal Execution, Limit Order Book, Reinforcement Learning, Agent-Based Simulation, Algorithmic Trading, Transaction Costs, Market Impact.
@@ I Introduction
In the intricate landscape of financial markets, optimizing the execution of significant positions remains a critical challenge for financial institutions such as banks, asset management firms, hedge funds, and prop shops.
@@
Research in market microstructure shows that large trades influence asset prices because the immediate depth of the market is limited [Bouchaud et al., 2009]; a single large order can exhaust all current buyers or sellers. This suggests that it is generally advantageous to split large orders into several smaller blocks. Furthermore, nuances in portfolio adjustment can cause adverse price changes, forcing traders to strike a delicate balance between quick trading and possible poor execution, or slower trading and exposure to unpredictable market fluctuations [Gueant, 2016].
Various techniques are used to minimize market impact, which generally consist of splitting the large order into smaller ones. Traditional approaches, typified by the Almgren-Chriss approach [Almgren and Chriss, 2001], rely on stochastic optimal control to maximize performance and are solved analytically. For example, [Almgren, 2012] explored the crafting of optimal execution strategies in the context of varying market liquidity and volatility over time. [Cartea et al., 2018] developed a high-frequency trading strategy, using superior speed for information processing and order placement while introducing a multifactor mutually exciting process to enable feedback effects on market orders and the shape of the limit order book. Meanwhile, [Giacinto et al., 2022] examined optimal liquidation within a market containing various diverse market makers with constrained inventory-holding and risk-bearing capacities.
However, the approach to solving the stochastic control problem is limited. Analytical solutions are rare and require stringent conditions on the problem’s modeling. Other approaches involve solving or approximating the Hamilton-Jacobi-Bellman (HJB) equation (see [Bertsekas, 2005]) or Quasi-Variational inequalities (see [Carmona and Delarue, 2002, Øksendal and Sulem, 2007]), or even addressing the problem through its probabilistic formulation using backward stochastic differential equations (BSDEs) [El Karoui and Quenez, 1997]. These methods are constrained and often rely on viscosity solutions to prove the existence and uniqueness and characterize the regularity of the value function [Crandall and Lions, 1992]. Furthermore, numerical schemes to approximate these equations, such as Monte-Carlo methods (see [Glasserman, 2004]) for BSDEs and other discretization methods for PDEs, struggle with high-dimensional problems (beyond three to four dimensions), significantly limiting the modeling capability.
Reinforcement learning techniques have emerged as viable solutions, exploring sequential decision problems without stringent modeling assumptions. These methodologies, applied in solving the optimal execution problem provide avenues when analytical solutions are elusive.
One of the challenges in studying this problem is that, in historical simulations, it is not possible to reproduce the effect of the trade without making assumptions about market impact. Therefore, we address this issue using the multiagent market simulator ABIDES [Byrd et al., 2020] to train and test our approach. This approach mitigates the challenges encountered when studying this problem, especially in historical simulations, where accurately reproducing trade effects without assuming market impact is difficult. To stabilize the profit and loss (P&L) dynamics, we impose penalties on trade-throughs—instances where at least one limit in the order book is depleted. Additionally, we implement a delayed penalty reward to ensure that agents do not retain any remaining inventory at the end of the execution period.
II The Limit Order Book
Figure 1: Representation of LOB
@@ II The Limit Order Book
A Limit Order Book (LOB) is a comprehensive record of current limit orders, where price changes occur discretely in increments known as the tick size (see Figure 1). Each order consists of a price and size, collectively contributing to the volume represented by $Q$. One of the key elements highlighted in the Limit Order Book is the concept of liquidity. In a perfectly liquid market, any quantity of a specific security can be immediately converted to cash. Measurement of liquidity involves metrics such as traded value / turnover, bid-ask spread, and trade-through intensity, as described in [Chevalier et al., 2023].
@@
  • Total depth: it corresponds to the cumulative sum of volume at multiple price levels ($d$) of the LOB:
[…]
@@
II-A Modelling the LOB
Traditional LOB models often overlook key market features, such as heavy tails and volatility clustering. The early works [Stigler, 1964] and [Garman, 1976] focused on market regulations and microstructure using simple simulations. Later models introduced different types of trades and mechanisms, such as order flow and herding behavior (see [Gould et al., 2013]). These new approaches leverage large datasets and focus on better replication of various stylized facts such as heavy tails and volatility clustering.
Stochastic Models
Stochastic models (also referred to as Zero Intelligence models) represent the dynamics of the LOB using probabilistic processes. These models often describe the arrival of orders and their execution using Poisson processes or other random mechanisms. Stochastic models are valuable for their analytical tractability and the ability to derive closed-form expressions for various market metrics. The seminal work [Cont et al., 2010] provides a comprehensive stochastic model for the LOB, capturing key features such as order flow and market depth, and a fast estimation method from market data. Other notable examples include [Smith et al., 2003, Huang et al., 2013, Tommaso Mariotti and Toscano, 2023].
Machine Learning Approaches
Machine learning techniques, particularly deep learning, have gained popularity in LOB modeling due to their ability to capture complex, nonlinear patterns in data. These methods do not require agent calibration, but instead learn simulated market behavior directly from historical data. Generative Adversarial Networks (GANS) [Cont et al., 2023, Li et al., 2020, Coletta et al., 2022a] are commonly used to model temporal dependencies in LOB data.
Agent-Based Models
Agent-Based Models (ABMs) simulate the interactions of autonomous agents, each following a set of rules or strategies. These models are particularly useful for capturing the heterogeneous and strategic behavior of market participants. ABMs offer insight into individual agent performance and the overall impact of their interactions. These models lie between zero-intelligence and perfect-rationality models, as they permit agent behaviors to be specified without the need for explicit rationality. Examples of ABMs for LOB modeling include [Alfi et al., 2009, Chakraborti et al., 2011, Hamill and Gilbert, 2015, Coletta et al., 2022b].
Why ABIDES?
After reviewing various approaches, we decided to use the ABIDES (Agent-Based Interactive Discrete Event Simulation) [Byrd et al., 2020] framework for our research. ABIDES is designed to simulate realistic financial markets with high fidelity. It supports the creation of diverse agent behaviors, including those of institutional and retail traders, market makers, and high-frequency traders. ABIDES allows for detailed modeling of LOB dynamics and the interactions between different market participants.
II-B ABIDES Configuration
@@
Let $(\Omega,\mathbb{F},\mathcal{F}=\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a complete filtered probability space endowed with right-continuous filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$. Going forward, whenever we mention equalities between random variables, we mean that they hold true almost surely under the probability measure $\mathbb{P}$. We may not explicitly indicate this notion ($\mathbb{P}$a.s.) in most cases. In the simulation, various agents, such as Exchange Agents, Adaptive Agents, Market Maker Agents, Value Agents, and Momentum Agents, develop their trading strategies based on an estimate of the fundamental value, which is defined by an Ornstein-Uhlenbeck (OU) process. The dynamics of the fundamental process process can be described by the following stochastic differential equation (SDE):
$dX_{t}=\theta(\mu-X_{t})\,dt+\sigma\,dW_{t}+J\,dN_{t},$ (1)
@@
  • $W_{t}$ is a standard Wiener process,
  • $N_{t}$ is a Poisson process with intensity $\lambda$, representing the arrival of major news events,
@@
The jump process $N$ is introduced to represent events related to extrinsic news of substantial nature that occur relatively infrequently but have the potential to significantly alter the consensus valuation of a stock. This allows us to obtain improvements of the simulation computation time while producing more realistic price time series.
We chose the RMSC-4 configuration, which is the reference configuration in ABIDES and was also used in [Amrouni et al., 2022] and [Shi and Cartlidge, 2023]. Several key settings are defined within the RMSC-4 configuration to realistically simulate market behaviors and participant interactions. Exchange agents manage order books and historical data streams, maintaining a depth of 10 levels and a history of 500 streams.
@@
Market Maker agents dynamically adjust their pricing strategies. They use an adaptive window size, set their order size to $0.025$ percent of the volume, maintain a price range within 10 ticks, and operate with a wake-up frequency of $1$ second to update their quotes. This setup uses a higher frequency compared to other standard configurations, making it more realistic and better adapted to today’s markets (e.g. [Nagy et al., 2023, Karpe et al., 2020]).
@@
These settings are designed to provide a comprehensive simulation of market dynamics, reflecting the diverse strategies and behaviors of market participants within a controlled environment.
Figure 2: Ask prices generated by ABIDES for different seeds.
III Optimal Execution Setting
The optimal execution problem consists in executing a trade of $X_{0}$ shares in a maximum amount of time $T$. In assessing the trade, it’s essential to consider its immediate transactional impact, temporary price fluctuations, and potential long-term effects on future prices due to lasting alterations in pricing dynamics. Additionally, a penalty could be considered if the agent fails to accomplish complete execution within the specified time period $[0,T]$. In what follows, we will consider an optimal liquidation problem without loss of generality.
In a discrete setting with a number of time-steps $N+1$, the execution strategy is a sequential decision process in which the trader decides the quantity to execute $x_{k}$ in each time step $t_{k}=k\frac{T}{N}$ for $k\in\{0,...,N\}$, where $t_{0}=0$ and $t_{N}=T$. A trading trajectory is defined as a list $\{x_{0},...,x_{N}\}$, where $x_{k}$ is the number of units held at time $t_{k}$. Clearly, $x_{0}=X_{0}$ and liquidation at time $T$ requires $x_{N}=0$, that is, $\sum_{k=0}^{N}x_{k}=X_{0}$. Let $P_{n}$ denote the average liquidation price for trade $x_{n}$ [^1] The execution price $P_{n}$ typically depends on both the current trade $x_{n}$ and all preceding trades. Additionally, we consider that $P_{n}$ encompasses all transaction costs.. The goal is to minimize the expected total liquidation cost [^2] This objective prioritizes expected value for a risk-neutral trader. It can be extended to include more risk factors. over the horizon $T$ :
$\min_{x\in\mathcal{A}}\mathbb{E}\left[\sum_{k=0}^{N}P_{k}x_{k}\right],$ (2)
$\mathcal{A}=\left\{\left\{x_{0},x_{1},\ldots,x_{N}\right\}\in\mathbb{R}^{N+1}_{+};\sum_{k=0}^{N}x_{k}=X_{0}\right\}.$ (3)
@@ III Optimal Execution Setting
where $P_{k}$ is the average at which each trade was executed through a market order sent at time $t_{k}$. In Almgren-Chriss terms [Almgren and Chriss, 2001], TWAP corresponds to a scenario with zero risk aversion, where the trader is indifferent to price volatility and only aims to minimize market impact by trading evenly.
IV Related Works
Initial studies on optimal execution are based on a stochastic framework in which the dynamics of the assets and execution costs are modeled through SDEs. The first formal analysis of the optimal execution problem is by [Bertsimas and Lo, 1998]. Under suitable conditions and using dynamic programming, they provide a closed-form formula as a solution to the optimal execution problem. Subsequently, an extension of the stochastic model was provided by introducing permanent and temporary effects due to the impacts of orders on the market and by inserting a risk-aversion parameter by [Almgren and Chriss, 2001]. The Almgren-Chriss model optimizes the trading strategy $x(t)$ over a time horizon $[0,T]$ to minimize total costs, and has a parameter to adjust the trade-off between minimizing market impact and execution costs.
[Huberman and Stanzl, 2005] extended the optimal execution framework by introducing more complex price impact functions and risk aversion parameters. Furthermore, the book by [Gueant, 2016] proposes an in-depth analysis and extension of these approaches. These methods, however, make strong assumptions on the underlying price movement or distributions.
Thanks to technological advances and data availability, [Nevmyvaka et al., 2006] applied RL for the first time to optimal execution strategies, with the objective of minimizing the implementation shortfall. Subsequently, [Hendricks and Wilcox, 2014] proposed to combine the Almgren-Chriss model with the Q-learning algorithm, to create a hybrid framework that executes a proportion of the AC trajectory based on the states in input. To address the high dimensions and complexity of the underlying dynamics, [Ning et al., 2018] used the Deep Q-Network (DQN) [Mnih et al., 2015], a combination of deep neural network and Q-learning, for optimal trade execution, addressing the curse of dimensionality issue faced by tabular Q-learning. [Lin and Beling, 2020], one of the most recent works, used PPO [Schulman et al., 2017a], to propose an optimal execution framework that can account for temporal correlations and make decisions based on LOB data using a sparse reward signal. These approaches suffer from the shortcomings of learning from historical data: the impossibility of reproducing realistically the impact of the orders of the agent.
This development has facilitated multiagent approaches [Balch et al., 2019] and led to the creation of ABIDES [Byrd et al., 2020]. [Karpe et al., 2020] was the first to utilize ABIDES to simulate a realistic trading environment and applied the DDQL algorithm (e.g. [Van Hasselt et al., 2016]) to learn optimal execution policies. Furthermore, [Nagy et al., 2023] formulated an execution strategy based on predictive signals of future price movements. While our research addresses the practical challenges associated with executing large orders, [Nagy et al., 2023] focuses on integrating and optimizing predictive trading signals using Q-learning methods. In contrast, [Karpe et al., 2020] highlights the dynamic and interactive nature of agent behavior and collective strategy optimization in trading environments.
V Reinforcement learning

### v2

@@ Abstract
This study investigates the development of an optimal execution strategy through reinforcement learning, aiming to determine the most effective approach for traders to buy and sell inventory within a finite time horizon. Our proposed model leverages input features derived from the current state of the limit order book and operates at a high frequency to maximize control. To simulate this environment and overcome the limitations associated with relying on historical data, we utilize the multi-agent market simulator ABIDES, which provides a diverse range of depth levels within the limit order book. We present a custom MDP formulation followed by the results of our methodology and benchmark the performance against standard execution strategies. Results show that the reinforcement learning agent outperforms standard strategies and offers a practical foundation for real-world trading applications.
@@ I Introduction
Research in market microstructure shows that large trades influence asset prices because the immediate depth of the market is limited (see [1]); a single large order can exhaust all current buyers or sellers. This suggests that it is generally advantageous to split large orders into several smaller blocks.
There is an extensive stream of literature on how to optimally split orders, which will be introduced in Section II, but the majority of this research restricts market behavior to Brownian motion with hard coded market impact assumptions and does not simulate the liquidity present in the markets, obtaining an execution trajectory which is a function of time.
Reinforcement learning (RL) is a very flexible machine learning paradigm which can be applied to learn data-driven execution policies, without having to make specific assumptions on market behavior. On the other hand, RL algorithms are very data hungry and require thousands of simulations to be trained, for this reason having a market simulator is essential.
One can simulate the market using stochastic processes and making assumptions on the market impact, but then the optimal approach is to use the traditional solutions which are optimal in that context [2]. Another solution can be to use historical data, as is done by several in the optimal execution with RL framework ([3, 4]), which gives access to LOB information, but in general enables only a simulation of temporary impact as the price must revert back to the historical dataset while in a realistic framework there is also temporary and transient impact (see Section III-B).
The approach considered in this paper is through the use of a multi agent market simulation, which enables the simulation of the entire LOB and takes into account the impact of the trades in the market. Examples of such simulators for LOB modeling include [6, 7, 8, 9]. This development has facilitated multi-agent approaches like [10] and led to the creation of ABIDES [11] which is the simulator considered in this work.
In this paper we present a framework based on RL, specifically the Double Q Network (DQN) [12] algorithm, to learn an optimal execution strategy. Training and testing occur on synthetic data generated through the multi-agent market simulator ABIDES.
The rest of the paper is organized as follows, Section II reviews the related works, Section III introduces the limit order book, market impact, the multi agent simulation and then the optimal execution problem setting. Section IV introduces RL and explains how the optimal execution problem can be solved with this technology. Section V presents our experimental results.
II Related Works
We will divide the literature on optimal execution into two major strands: the classical mathematical finance framework and the more recent RL approaches. Within the mathematical finance tradition, [13] first established a model where prices follow Brownian motion with linear permanent and temporary market impact, deriving the Time-Weighted Average Price (TWAP) as the optimal solution. This was extended by [14], who augmented the objective function with the minimization of the variance of the expected shortfall, yielding an analytical solution that incorporates risk aversion—resulting in more front-loaded execution profiles. [15] further refined the model by introducing transient impact that decays over time, producing a U-shaped execution trajectory with higher trading intensity at the beginning and end of the period. Building on this, Busetti and Lillo [16] introduced risk aversion into the transient impact model, confirming the U-shaped profile but with heavier emphasis on early execution. [17] expanded this line of work by testing alternative objective functions, showing that inventory penalties act analogously to risk aversion and exploring terminal penalties for unfinished execution.
A common feature across these studies is their focus on what we may term strategic execution: the optimal trading trajectory depends solely on deterministic variables such as time, permanent impact, and risk aversion, without regard to real-time market microstructure or limit order book dynamics. The execution period lasts a few hours, with usually five minute time-steps. The other type of execution which can be analyzed is the tactical execution, where the focus is to optimize the trade decisions within each period. In this case, LOB information is fundamental.
Reinforcement learning methods allow for such tactical execution strategies, thus we will now analyze them by dividing into the two categories. The strategic RL approaches have in general the objective of proving RL can find the optimal solutions found by traditional methods. In this category are the following papers: [3, 4, 2], which typically operate over long time windows and use algorithms such as Q-learning, DQN, or PPO to learn execution policies from historical or simulated limit order book data.
Conversely, tactical RL approaches deal with shorter time horizons and more granular decision-making. Early work by Nevmyvaka et al. [18] pioneered this direction, optimizing execution over five-minute intervals. Subsequent contributions [19, 20, 21], investigated higher frequency control problems such as deciding between market and limit orders, stopping execution under uncertainty, or leveraging predictive signals to guide trades.
In this work, we aim to unify these two traditionally separate layers of the execution problem. By training a reinforcement learning agent to operate over a 30-minute window with one-second control intervals, we enable the simultaneous learning of both strategic scheduling and tactical decision-making within a single framework, bridging the gap between high-level trade planning and low-level order placement in real time.
The most similar to our approach is [22], as they use ABIDES as the market simulator and they have a high (30 second) control frequency to learn both a strategic and tactical execution. Nonetheless, their experimental results are unsatisfactory as they state that their RL agent solution converges to the TWAP. Instead, in this work we have run an extensive experimental campaign obtaining results superior to the TWAP and other baseline strategies.
III Optimal Execution Setting
III-A Limit Order Book
@@
A limit order book is a comprehensive record of current limit orders, where price changes occur discretely in increments known as the tick size (see Figure 1). Each order consists of a price and size, collectively contributing to the volume represented by $Q$. One of the key elements highlighted in the LOB is the concept of liquidity.
Figure 1: Representation of the Limit Order Book
In a perfectly liquid market, any quantity of a specific security can be immediately converted to cash. Measurement of liquidity involves metrics such as traded value/turnover, bid-ask spread, and trade-through intensity, as described in [23].
@@
  • Total depth: it corresponds to the cumulative sum of volume at multiple price levels $d$ of the LOB:
    $TD_{{h}}^{k}=$ $\sum_{j=1}^{k}Q_{{h}}^{j}\;\text{for}\;k\in\{1,...,d\},\;{h}\in\{\text{bid, ask}\},$
@@
III-B Market impact
Market impact refers to the effect that trades have on asset prices and is a key consideration in the design of optimal execution strategies. The literature typically distinguishes between three primary forms of market impact:
  • Temporary Impact: This refers to the immediate but short-lived distortion in price caused by a trade. It is generally attributed to the consumption of available liquidity at the best bid or ask and tends to vanish quickly as new orders replenish the book or the market reverts. Temporary impact is typically modeled as a cost that affects only the execution price of the current trade.
  • Permanent Impact: Permanent impact captures the long-lasting effect that a trade has on the asset price. It is often interpreted as the market incorporating the information revealed by the trade, such as the presence of an informed trader or a shift in supply and demand. Permanent impact is usually modeled as a linear function of cumulative volume and affects the price trajectory permanently.
  • Transient Impact: Transient impact lies between the two extremes above. It reflects a price impact that decays gradually over time rather than disappearing instantaneously (as with temporary impact) or persisting indefinitely (as with permanent impact). This type of impact is useful for modeling situations where market participants adjust their behavior in response to past trades, leading to delayed and decaying price effects. Transient impact is often captured through decaying kernels or memory functions in execution models.
Figure 2: Stylized representation of average permanent vs. temporary price impact (see [24]).
ABIDES reproduces realistic market impact because price responses emerge endogenously from the interaction of heterogeneous agents rather than from imposed functional forms. Large orders deplete liquidity at the best quotes, prompting market makers and other agents to adjust their behavior. This naturally generates short-lived distortions resembling temporary impact as well as longer-lasting adjustments consistent with permanent and transient impact. Thus, the dynamics of impact arise from microstructural interactions, providing a realistic environment for testing execution strategies.
III-C Configuration of Multi-Agent Simulations
@@
Let $(\Omega,\mathbb{F},\mathcal{F}=\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a complete filtered probability space endowed with right-continuous filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$. We assume that the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ supports a standard $\mathbb{P}$-Brownian motion $W$ and a random Poisson process $N$ with intensity $\lambda$. In the simulation, various agents, such as Exchange Agents, Adaptive Agents, Market Maker Agents, Value Agents, and Momentum Agents, develop their trading strategies based on an estimate of the fundamental value, which is defined by an Ornstein-Uhlenbeck (OU) process. The dynamics of the fundamental process process can be described by the following stochastic differential equation (SDE):
$\mathrm{d}X_{t}=\theta(\mu-X_{t})\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t}+J\,\mathrm{d}N_{t},$
@@
The jump process $N$ is introduced to represent the impact of the arrival of major news events on the fundamental value of the traded asset. This allows us to obtain improvements of the simulation computation time while producing more realistic price time series.
We chose the RMSC-4 configuration, which is the reference configuration in ABIDES and was also used in [25] and [Shi and Cartlidge(2023)]. Several key settings are defined within the RMSC-4 configuration to realistically simulate market behaviors and participant interactions. Exchange agents manage order books and historical data streams, maintaining a depth of 10 levels and a history of 500 streams.
@@
Market Maker agents dynamically adjust their pricing strategies. They use an adaptive window size, set their order size to $0.025$ percent of the volume, maintain a price range within 10 ticks, and operate with a wake-up frequency of $1$ second to update their quotes. This setup uses a higher frequency compared to other standard configurations, making it more realistic and better adapted to today’s markets (e.g. [21, 22]).
@@
These settings are designed to provide a comprehensive simulation of market dynamics, reflecting the diverse strategies and behaviors of market participants within a controlled environment, as well as simulate realistic market impact.
Figure 3: Sample paths of ask prices generated by ABIDES for different seeds.
III-D Problem Formulation
The optimal execution problem consists in executing a trade of $X_{0}$ shares over a time interval $[0,T]$. In assessing the trade, it’s essential to consider its immediate transactional impact, temporary price fluctuations, and potential long-term effects on future prices due to lasting alterations in pricing dynamics. Additionally, a penalty could be considered if the agent fails to accomplish complete execution on time.
In a discrete-time setting with N+1N{+}1 decision points, the purchasing strategy (analogously for selling) is formulated as a sequential decision process, in which the trader determines the quantity $x_{k}$ to buy at each time step $t_{k}=kT/N$, for $k\in\{0,\ldots,N\}$, with $t_{0}=0$ and $t_{N}=T$. The cumulative purchases form a trajectory $\{x_{0},\ldots,x_{N}\}$, where $x_{k}$ denotes the cumulative number of shares acquired by time $t_{k}$. By construction, $x_{0}=0$, and full completion requires $x_{N}=X_{0}$ at the terminal time. Let $P_{k}$ denote the average execution price for trade $\Delta x_{n}=x_{k}-x_{k-1}$. The goal is to minimize the expected total execution cost for a risk-neutral trader over a finite time horizon $T$ :
$\min_{x\in\mathcal{A}}\mathbb{E}\left[\sum_{k=0}^{N}P_{k}\Delta x_{k}\right],$ (1)
$\mathcal{A}=\left\{\left\{x_{0},x_{1},\ldots,x_{N}\right\}\in\mathbb{R}^{N+1}_{+};\sum_{k=0}^{N}|\Delta x_{k}|=X_{0}\right\}.$ (2)
This cost functional corresponds to the Implementation Shortfall (IS) introduced by [26], which measures the excess cost of execution relative to an instantaneous purchase at the initial market price $P_{0}$. Formally, for a purchasing strategy, the IS is
$\text{IS}=\sum_{k=0}^{N}P_{k}\,\Delta x_{k}-X_{0}P_{0}.$
Initial studies on optimal execution are based on a stochastic framework in which the dynamics of the assets and execution costs are modeled through SDEs. The first formal analysis of the optimal execution problem is by [13]. Under suitable conditions and using dynamic programming, they provide a closed-form formula as a solution to the optimal execution problem. Subsequently, an extension of the stochastic model was provided by introducing permanent and temporary effects due to the impacts of orders on the market and by inserting a risk-aversion parameter by [14]. The Almgren-Chriss model optimizes the trading strategy $x(t)$ over a time horizon $[0,T]$ to minimize total costs, and has a parameter to adjust the trade-off between minimizing market impact and execution costs.
@@ III-D Problem Formulation
where $P_{k}$ is the average at which each trade was executed through a market order sent at time $t_{k}$. In Almgren-Chriss terms [14], TWAP corresponds to a scenario with zero risk aversion, where the trader is indifferent to price volatility and only aims to minimize market impact by trading evenly.
IV Reinforcement Learning